# A Bayesian Estimator of Sample Size

### v2

@@ 4.1 Correlation between sample size, evidence, and confidence
Figure 1: The line plots of (left) confidence vs. evidence when sample size is fixed to be $n=10,20,\text{ and }30$, and (right) confidence vs. sample size when $e=0.1$. The result assumes binary outcome for two-arm trial, with $\theta^{*}=0.05$.
@@
We again compute confidence using equation (3), varying the sample size $n$ while keeping the evidence $e$ constant. Figure 1(b) demonstrates the positive relationship between sample size and confidence, when $e=0.1$ and $\theta^{*}=0.05$. This is expected when evidence is supportive of the alternative hypothesis since the larger the sample size , the larger the posterior probability of $H_{1}$.
@@
We next explore the correlation between evidence and sample size, maintaining a fixed confidence level at $\text{Pr}(H=H_{1}|e,n)=0.6.$ For this demonstration, we calculate confidence using (3) across a range of evidence and sample sizes. For example, with $n=10$ patients per arm, the potential range of evidence $e$ reflecting the difference in percent responders between the treatment and control arms - spans from -1 to 1, with increments of 0.1.
Given the discrete nature of the outcomes, it’s not always possible to find an evidence level for each sample size $n$ that exactly matches $\text{Pr}(H=H_{1}|e,n)=0.6$. Thus, we document the smallest $e$ for each $n$ where $\text{Pr}(H=H_{1}|e,n)\geq 0.6$. In this case, for $\theta^{*}=0.05$, we report $n(e)$, the different sample sizes $n$ with their corresponding evidence $e$: $50(0.080)$, $100(0.070)$, $150(0.067)$, $200(0.065)$, and $1000(0.057)$. These values clearly show that as sample size increases, evidence needed to maintain the confidence at 0.6 decreases, approaching $\theta^{*}=0.05$.
@@ 4.2 Coherence between BESS and Bayesian Inference
After the trial is conducted and data observed, a Bayesian analysis of the observed data using posterior probability will be coherent with the BESS statement. To explain this, we first denote $\bm{y}^{\star}_{n}$ the observed data after the trial is completed, $e^{\star}$ the observed evidence, and $c^{\star}=\text{Pr}(H_{1}|\bm{y}^{\star}_{n})$ the posterior probability of $H_{1}$ conditional on the observed data $\bm{y}^{\star}_{n}.$ Then the coherence between BESS and Bayesian inference means:
If $e^{\star}>e$, then $c^{\star}>c.$
@@
To verify the coherence property numerically, we perform a simple simulation. In Appendix Table A.1 we set up the true parameters for data simulation for each type of clinical trial data. For example, for two-arm binary data, we assume the minimum treatment effect $\theta^{*}=0.05$, evidence $e=0.1$, confidence $c=0.8$. Using a prior distribution $\theta_{j}\sim\text{Beta}(0.5,0.5)$ BESS leads to a sample size of 150. We then repeatedly generate trial data assuming different true values of $\theta_{1}$ and $\theta_{0}$ in Table A.1, and report the inference results of $e^{\star}$ and $c^{\star}$. Figure 2 summarizes the simulation results. In all three subplots (a)-(c), when $e^{\star}$ is larger than $e$ represented by the black vertical line, $c^{\star}$ is larger than $c$, the red horizontal line.
@@ 5.2 Simulation Result
Table 2 presents the simulation results. They are both surprising and reassuring that BESS and the standard SSE produce similar sample sizes across a variety of settings. It is surprising since the two approaches are based on different statistical metrics. For BESS, it’s aiming to balance between the anticipated evidence in the observed data and the confidence expressed as posterior probability; for the standard approach, it’s trading off among the type I/II error rates and assumed true parameter values. The results are reassuring since despite using different metrics, when matching the type I/II error rates, both approaches produce highly similar sample size estimates.
Several trends are worth noting in Table 2. First, increasing confidence cutoff $c$ leads to lower type I error rate for fixed evidence $e.$ This is because increase in cutoff $c$ makes it harder to reject the null for BESS under (7), and hence fewer simulated trials under $H_{0}$ will be rejected, i.e., decrease in the type I error rate. Second, note that the true effect size under $H_{1}$ equals $\theta\equiv\theta_{1}-\theta_{0}=0.4-0.25=0.15.$ We observe that 1) power increases when the confidence cutoff $c$ increases and evidence $e$ is less than the true effect size, i.e., $e=0.1,$ 2) power stays the same if evidence $e$ equals the true effect size, and 3) power decreases when confidence increases and $e$ is greater than the true effect size. This complicated trend demonstrates an interaction between $c$ and power conditional on whether $e$ is greater than, equal to, or less than the true effect size $\theta$. To understand this, first recall in Section 4.2 we used notation $e^{\star}$ to denote the observed evidence after the trial is completed and data observed. When $c$ increases, the BESS-estimated sample size also increases (see Section 4.1). Therefore, the observed $e^{\star}$ will be closer to $\theta$ due to large number theory. Consequently, more simulated trials under the alternative will see $e^{\star}$ close to $\theta$. If the assumed evidence $e$ is smaller than $\theta$, more likely it will be smaller than $e^{\star}$ as well. This means that the posterior probability $\text{Pr}(H_{1}|\bm{y}_{n})$ will be higher (since there is stronger evidence supporting $H_{1}$), and hence more rejections, i.e., higher power. Therefore, when $e$ is smaller than $\theta$, a larger $c$ leads to higher power. Same logic applies to the case when $e$ is larger than $\theta,$ in which a larger $c$ leads to lower power. Lastly, when $e=\theta$, increasing sample size (as a result of increasing $c$) makes $e^{\star}$ approach $e$, and therefore there is no obvious impact on the power.
@@
Finally, from a true Bayesian perspective, BESS is concerned about the trial at hand, rather than hypothetical trials generated from the null or alternative (i.e., type I/II error rates). Therefore, while Table 2 illustrates a connection between BESS and standard SSE, it does not imply that BESS needs to be calibrated based on type I/II error rates in practice. On the contrary, BESS focuses on the probability of making a right decision given the observed data, which can be measured by the FPR and FNR. In Table 2 the reported FPR and FNR for BESS and standard SSE are the same since 1) the prevalence of trials under the $H_{0}$ and $H_{1}$ is 50% and 2) the standard SSE is oracle since it assumes the true $\theta_{0}$ and $\theta_{1}$ values in its computation. In Appendix Table A.2 we show that when the $\theta_{0}$ and $\theta_{1}$ are mis-specified in the standard SSE, the estimated sample sizes may be too large or too small, leading to over- or under-power, and deflated or inflated FPR/FNR’s.
Evidence Confidence type I power BESS Standard SSE
$e$ $c$ error rate $\alpha$ $1-\beta$ $n$ FPR FNR $n$ FPR FNR
0.10 0.7 0.31 0.76 60 0.29 0.26 62 0.29 0.26
0.8 0.20 0.84 150 0.19 0.17 145 0.19 0.17
0.9 0.10 0.94 340 0.10 0.07 344 0.10 0.07
0.15 0.7 0.29 0.56 20 0.34 0.38 22 0.34 0.38
0.8 0.21 0.56 40 0.27 0.36 40 0.27 0.36
0.9 0.10 0.56 87 0.16 0.33 88 0.16 0.33
0.20 0.7 0.44 0.57 5 0.43 0.43 5 0.43 0.43
0.8 0.24 0.46 15 0.34 0.41 16 0.34 0.41
0.9 0.11 0.38 35 0.23 0.41 36 0.23 0.41
@@
We demonstrate the sensitivity of incorporating prior information through simulation, assuming there exists prior data of $n_{0}$ patients per arm. The simulation details are presented in Appendix A.6, and the average sample size from the 1,000 simulated trials is 31.42 with a standard deviation of 28. Recall with a Beta(0,0)00(0,0)( 0 , 0 ) prior the BESS sample size was 40 in Table 2 for $e=0.15$ and $c=0.8$. The results show that BESS is able to properly borrow prior information for sample size estimation. In practice, one may use different informative priors, e.g, power prior (Ibrahim and Chen, 2000) or commensurate prior (Hobbs et al., 2011), for information borrowing. We leave these topics for future research.
@@ 6.1 Fixed Sample Size
BESS provides the following sample size statement: Assuming evidence $e=0$, a sample size of 57 patients per arm is needed to declare with 70% confidence that the response rate of the higher dose is no higher than the lower dose by 0.05. Larger or smaller sample size can be achieved by calibrating $e$ and $c$.
@@
Assume 20 patients per dose are randomized and denote $\bar{y}_{L}$ and $\bar{y}_{H}$ the observed response rates for the lower and higher doses, respectively. We compute the confidence, $\text{Pr}(\theta_{H}-\theta_{L}<\theta^{*}|\bar{y}_{L},\bar{y}_{H},n)$, for various values of $(\bar{y}_{L}-\bar{y}_{H})$ and $\theta^{*}=0.05,$ shown in Table 3. For example, if one observes the response rate of the lower dose is the same as the higher dose, i.e., $(\bar{y}_{L}-\bar{y}_{H})=0$, with 20 patients per dose one gets about 62.49% confidence to declare the lower dose response rate is within the non-inferiority margin $0.05$ of the higher dose.
@@ 1. BESS SSR
The first design is BESS with sample size re-estimation (SSR). BESS SSR estimates a sample size $n$ for the entire trial first using input of evidence $e$ and confidence $c$. Then when $n/2$ patients are enrolled at each dose, an interim analysis is performed to allow trial stopping or SSR if the trial is not stopped. Denoting the interim patients outcome as $\bm{y}_{n/2}$, the trial is stopped early if $\text{Pr}(H_{1}|\bm{y}_{n/2})\geq c$ or $\text{Pr}(H_{1}|\bm{y}_{n/2})\leq c^{*}$, where $c$ close to 1 and $c^{*}$ close to 0 are probability thresholds for early stopping due to success or failure, respectively. If neither condition is met, the trial proceeds with an SSR as follows.
@@ 3. Standard SSE
The standard SSE estimates a sample size $n$ based on a $z$-test for (9), with a significance level $\alpha$, and a desired power $(1-\beta).$ The null hypothesis is rejected if $1-\Phi(z)\leq\alpha$. Formula (6) is used to compute $n.$
@@ 4. Standard SSE with interim
$\theta_{L}|\theta_{H},H\sim\left\{\begin{matrix}\text{Unif}(0,\theta_{H}-\theta^{*})&H=0\\ \text{Unif}(\theta_{H}-\theta^{*},0.28)I(\theta_{H}-\theta^{*}\leq 0.28)+\delta_{\theta_{H}}I(\theta_{H}-\theta^{*}>0.28)&H=1\end{matrix}\right.$
This setting ensures $E(\theta_{L})=E(\theta_{H})=0.335$ when $H=1$. In scenario 2, we assume $\theta_{H}=0.335$ and $\theta_{L}=0.265$ under $H_{0}$ or $\theta_{L}=0.335$ under $H_{1}$, i.e., they are fixed. The two scenarios reflect two different practical considerations. Scenario 1 aims to assess the performance of the four designs assuming they are applied to different programs and trials in which the true responses of the drugs and doses are different. Scenario 2 is a classical Frequentist setting to assess the Type I/II error rates of a design assuming true response rates are fixed. Finally, given $\theta_{L}$ and $\theta_{H}$, for each trial we simulate patients outcome data based on Binomial distributions and the corresponding designs.
We assume $q=0.5$. For Designs 1 & 2, we let $e=0.02,$ $c=0.7$ and $c^{*}=0.3$. For designs 3 & 4, we set $\alpha=0.3$ and $\beta=0.3$, so that BESS and standard SSE produce the same sample size for the trial in the beginning, which is $n=100.$ Lastly, we assume $(\theta_{L},\theta_{H})\sim\text{Beta}(0.05,0.05)\text{Beta}(0.05,0.05)I(\theta\in H)$, where $\text{Beta}(0.05,0.05)$ is chosen to reduce the prior effective sample size. A total of 2,000 simulated trials are generated, 1,000 each under $H_{0}$ or $H_{1}$, for each design in each scenario.
@@ Results
We first present the operating characteristics of the four designs by evaluating their type I and II error rates across each scenario. To facilitate the comparison of designs, motivated by Kim and Choi (2021) we consider a combined error rate defined as
@@
where the weight $k\in[0,\infty)$ is a pre-determined factor that quantifies the relative weight between the type I and type II error rates. We let $k\in\{0.5,1,1.5\}$. A lower CER is more desirable. Top two rows in Figure 3 reports the results for the four designs. Across all designs, as the sample size increases, the rate at which CER declines, indicating potential diminishing return when sample size gets larger. In other words, the gain in reduction of error rates may lessen when sample size continues to increase. Considering the substantial costs associated with patient enrollment, these findings suggest finding a “sweet spot” of sample size that achieves a desirable tradeoff between cost and statistical properties. Comparing across designs, we find that except for the Standard SSE design, the performance of the other designs are similar for the same average sample size. Notably, the BESS SSR Cap design (design 2) seems to be in general the best across most cases as it shows slightly lower CER for fixed sample size. In Scenario 2 when $k=0.5$, the standard SSE is the winning design. In contrast, it is the losing design in the same scenario but with $k=1.5.$ This seems to suggest the standard SSE is a better design if the Frequentist type I error rate is of important consideration in design evaluation. However, in early-phase dose optimization trials, type I error rate is not the primary concern. In fact, one may argue it is of the least concern. For example, one would be much more concerned if a “GO” decision that recommend a wrong dose to further clinical development, or a “No Go” decision that fails to recommend a promising dose. These are measured by the FPR and FNR in the bottom two rows of Figure 3.

### v3

@@ 4.1 Correlation between sample size, evidence, and confidence
Figure 1: The line plots of (left) confidence vs. evidence when sample size is fixed to be $n=10,20,\text{ and }30$, and (right) confidence vs. sample size when $e=0.1$. The result assumes binary outcome for a two-arm trial, with $\theta^{*}=0.05$.
@@
We again compute confidence using Equation (3), varying the sample size $n$ while keeping the evidence $e$ fixed. Figure 1(b) illustrates the positive relationship between sample size and confidence, when $e=0.1$ and $\theta^{*}=0.05$. This is expected when the evidence is supportive of the alternative hypothesis since the larger the sample size, the larger the posterior probability $H_{1}$.
@@
We next explore the correlation between evidence and sample size, maintaining a fixed confidence level at $\text{Pr}(H=H_{1}|e,n)=0.7.$ For this demonstration, we calculate confidence using (3) across a range of evidence and sample sizes. For example, with $n=10$ patients per arm, the potential range of evidence $e$ reflecting the difference in percent responders between the treatment and control arms - spans from -1 to 1, with increments of 0.1.
Given the discrete nature of the outcomes, it’s not always possible to find an evidence level for each sample size $n$ that exactly matches $\text{Pr}(H=H_{1}|e,n)=0.7$. Thus, we document the smallest $e$ for each $n$ where $\text{Pr}(H=H_{1}|e,n)\geq 0.7$. In this case, for $\theta^{*}=0.05$, we report $n(e)$, the different sample sizes $n$ with their corresponding evidence $e$: $50(0.100)$, $100(0.080)$, $150(0.073)$, $200(0.070)$, and $1000(0.059)$. These values clearly show that as sample size increases, evidence needed to maintain the confidence at 0.7 decreases, approaching $\theta^{*}=0.05$.
@@ 4.2 Coherence between BESS and Bayesian Inference
After the trial is conducted and data observed, a Bayesian analysis of the observed data using posterior probability will be coherent with the BESS statement if the prior used for sample size estimation and statistical inference is the same. To explain this, we first denote $\bm{y}^{\star}_{n}$ the observed data after the trial is completed, $e^{\star}$ the observed evidence, and $c^{\star}=\text{Pr}(H_{1}|\bm{y}^{\star}_{n})$ the posterior probability of $H_{1}$ conditional on the observed data $\bm{y}^{\star}_{n}.$ Then the coherence between BESS and Bayesian inference means:
If $e^{\star}\geq e$, then $c^{\star}\geq c.$
@@
To verify the coherence property numerically, we perform a simple simulation. In Appendix Table A.1 we set up the true parameters for data simulation for each type of clinical trial data. For example, for two-arm binary data, we assume the minimum treatment effect $\theta^{*}=0.05$, evidence $e=0.1$, confidence $c=0.8$. Using a prior distribution $\theta_{j}\sim\text{Beta}(0.5,0.5)$ BESS leads to a sample size of 110. We then repeatedly generate trial data assuming different true values of $\theta_{1}$ and $\theta_{0}$ in Table A.1, and report the inference results of $e^{\star}$ and $c^{\star}$. Appendix Figure A.1 (one-arm) and Figure 2 (two-arm) summarizes the simulation results. In all three subplots (a)-(c), when $e^{\star}$ is larger than $e$ represented by the black vertical line, $c^{\star}$ is larger than $c$, the red horizontal line.
@@ 5.1 Simulation Setup
Evidence Confidence Type I Power BESS Standard SSE
$e$ $c$ error rate $\alpha$ $1-\beta$ $n$ FPR FNR $n$ FPR FNR
0.10 0.7 0.36 0.73 40 0.33 0.29 42 0.33 0.29
0.8 0.23 0.82 120 0.22 0.19 117 0.22 0.19
0.9 0.12 0.92 290 0.12 0.08 284 0.12 0.08
0.15 0.7 0.35 0.56 14 0.38 0.39 14 0.38 0.39
0.8 0.23 0.56 34 0.29 0.36 36 0.29 0.36
0.9 0.12 0.56 74 0.18 0.34 74 0.18 0.34
0.20 0.7 0.45 0.59 5 0.43 0.43 6 0.43 0.43
0.8 0.28 0.51 15 0.35 0.41 16 0.35 0.41
0.9 0.19 0.48 35 0.28 0.39 39 0.28 0.39
@@ 5.2 Simulation Result
Table 2 presents the simulation results. They are both surprising and reassuring that BESS and the standard SSE produce similar sample sizes across a variety of settings. It is surprising since the two approaches are based on different statistical metrics. For BESS, it’s aiming to balance between the anticipated evidence in the observed data and the confidence expressed as posterior probability; for the standard approach, it’s trading off among the type I/II error rates and assumed true parameter values. The results are reassuring since despite using different metrics, when matching the type I/II error rates, both approaches produce highly similar sample size estimates. For other outcome and trial types, similar result can be find in Appendix Table A.2
Several trends are worth noting in Table 2. First, increasing confidence cutoff $c$ leads to lower type I error rate for fixed evidence $e.$ This is because increase in cutoff $c$ makes it harder to reject the null for BESS under (7), and hence fewer simulated trials under $H_{0}$ will be rejected, i.e., decrease in the type I error rate. Second, note that the true effect size under $H_{1}$ equals $\theta\equiv\theta_{1}-\theta_{0}=0.4-0.25=0.15.$ We observe that 1) power increases when the confidence cutoff $c$ increases and evidence $e$ is less than the true effect size, i.e., $e=0.1,$ 2) power stays the same if evidence $e$ equals the true effect size, and 3) power decreases when confidence increases and $e$ is greater than the true effect size. This complicated trend demonstrates an interaction between $c$ and power conditional on whether $e$ is greater than, equal to, or less than the true effect size $\theta$. To understand this, first recall in Section 4.2 we used notation $e^{\star}$ to denote the observed evidence after the trial is completed and data observed. When $c$ increases, the BESS-estimated sample size also increases (see Section 4.1). Therefore, the observed $e^{\star}$ will be closer to $\theta$ due to large number theory. Consequently, more simulated trials under the alternative will see $e^{\star}$ close to $\theta$. If the assumed evidence $e$ is smaller than $\theta$, more likely it will be smaller than $e^{\star}$ as well. This means that the posterior probability $\text{Pr}(H_{1}|\bm{y}_{n})$ will be higher (since there is stronger evidence supporting $H_{1}$), and hence more rejections, i.e., higher power. Therefore, when $e$ is smaller than $\theta$, a larger $c$ leads to higher power. Same logic applies to the case when $e$ is larger than $\theta,$ in which case a larger $c$ leads to lower power. Lastly, when $e=\theta$, increasing sample size (as a result of increasing $c$) makes $e^{\star}$ approach $e$, and therefore there is no obvious impact on the power.
@@
Finally, from a Bayesian perspective, BESS is concerned about the trial at hand, rather than hypothetical trials generated from the null or alternative (i.e., type I/II error rates). Therefore, while Table 2 illustrates a connection between BESS and standard SSE, it does not imply that BESS needs to be calibrated based on type I/II error rates in practice. On the contrary, BESS focuses on the probability of making a right decision given the observed data, which can be measured by the FPR and FNR. In Table 2 the reported FPR and FNR for BESS and standard SSE are the same since 1) the prevalence of trials under the $H_{0}$ and $H_{1}$ is 50% and 2) the standard SSE is oracle since it assumes the true $\theta_{0}$ and $\theta_{1}$ values in its computation. In Appendix Table A.3 we show that when the $\theta_{0}$ and $\theta_{1}$ are mis-specified in the standard SSE, the estimated sample sizes may be too large or too small, leading to over- or under-power, and deflated or inflated FPR/FNR’s.
@@
We demonstrate the sensitivity of incorporating prior information through simulation, assuming there exists prior data of $n_{0}$ patients per arm. The simulation details are presented in Appendix A.6, and the average sample size from the 1,000 simulated trials is 23.63 with a standard deviation of 9.22. Recall with a Beta(0,0)00(0,0)( 0 , 0 ) prior the BESS sample size was 34 in Table 2 for $e=0.15$ and $c=0.8$. The results show that BESS is able to properly borrow prior information for sample size estimation. In practice, one may use different informative priors, e.g, power prior (Ibrahim and Chen, 2000) or commensurate prior (Hobbs et al., 2011), for information borrowing. We leave these topics for future research.
@@ 6.1 Fixed Sample Size
BESS provides the following sample size statement: Assuming evidence $e=0$, a sample size of 95 patients per arm is needed to declare with 70% confidence that the response rate of the higher dose is no higher than the lower dose by 0.05. Larger or smaller sample size can be achieved by calibrating $e$ and $c$.
@@
Assume 20 patients per dose are randomized and denote $\bar{y}_{L}$ and $\bar{y}_{H}$ the observed response rates for the lower and higher doses, respectively. We compute the confidence, $\text{Pr}(\theta_{H}-\theta_{L}<\theta^{*}|\bar{y}_{L},\bar{y}_{H},n)$, for various values of $(\bar{y}_{L}-\bar{y}_{H})$ and $\theta^{*}=0.05,$ shown in Table 3. For example, if one observes the response rate of the lower dose is the same as the higher dose, i.e., $(\bar{y}_{L}-\bar{y}_{H})=0$, with 20 patients per dose one gets about 57.46% confidence to declare the lower dose response rate is within the non-inferiority margin $0.05$ of the higher dose.
@@ 1. BESS SSR
The first design is BESS with sample size re-estimation (SSR). BESS SSR estimates a sample size $n$ for the entire trial first using input of evidence $e$ and confidence $c$. Then when $n/2$ patients are enrolled at each dose, an interim analysis is performed to allow trial stopping or SSR if the trial is not stopped. Denoting the interim data as $\bm{y}_{n/2}$, the trial is stopped early if $\text{Pr}(H_{1}|\bm{y}_{n/2})\geq c$ or $\text{Pr}(H_{1}|\bm{y}_{n/2})\leq c^{*}$, where $c$ close to 1 and $c^{*}$ close to 0 are probability thresholds for early stopping due to success or failure, respectively. If neither condition is met, the trial proceeds with an SSR as follows.
@@ 3. Standard SSE
The standard SSE estimates a sample size $n$ based on a $z$-test for (9), with a significance level $\alpha$, and a desired power $(1-\beta).$ The null hypothesis is rejected if $1-\Phi(z)\leq\alpha$ using the trial data of $n$ patients.Formula (6) is used to compute $n.$
@@ 4. Standard SSE with interim
$\theta_{L}|\theta_{H},H\sim\left\{\begin{matrix}\text{Unif}(0,\theta_{H}-\theta^{*})&H=0\;(H_{0})\\ \text{Unif}(\theta_{H}-\theta^{*},0.28)I(\theta_{H}-\theta^{*}\leq 0.28)+\delta_{\theta_{H}}I(\theta_{H}-\theta^{*}>0.28)&H=1\;(H_{1})\end{matrix}\right.$
This setting ensures $E(\theta_{L})<E(\theta_{H})-\theta^{*}$ under $H_{0}$, and $E(\theta_{L})=E(\theta_{H})=0.335$ under $H_{1}$. In scenario 2, we assume $\theta_{H}=0.335$ and $\theta_{L}=0.265$ under $H_{0}$ or $\theta_{L}=0.335$ under $H_{1}$, i.e., they are fixed. The two scenarios reflect two different practical considerations. Scenario 1 aims to assess the performance of the four designs assuming they are applied to different drug programs and trials in which the true responses of the drugs and doses are different. Scenario 2 is a classical Frequentist setting to assess the type I/II error rates of a design assuming true response rates are fixed. Finally, given $\theta_{L}$ and $\theta_{H}$, for each trial we simulate patients outcome data based on Binomial distributions and the corresponding designs.
We assume $q=0.5$. For Designs 1 & 2, we let $e=-0.02,$ $c=0.7$ and $c^{*}=0.3$. For designs 3 & 4, we set $\alpha=0.3$ and $\beta=0.3$, so that BESS and standard SSE produce the same sample size for the trial in the beginning, which is $n=100.$ Lastly, we assume $(\theta_{L},\theta_{H})\sim\text{Beta}(0.05,0.05)\cdot\text{Beta}(0.05,0.05)I(\theta\in H)$, where $\text{Beta}(0.05,0.05)$ is chosen to reduce the prior effective sample size. A total of 2,000 simulated trials are generated, 1,000 each under $H_{0}$ or $H_{1}$, for each design in each scenario.
@@ Results
We first present the operating characteristics of the four designs by evaluating their type I and II error rates (See Appendix A.7 for detail) across each scenario. To facilitate the comparison of designs, motivated by Kim and Choi (2021) we consider a combined error rate defined as
@@
where the weight $k\in[0,\infty)$ is a pre-determined factor that quantifies the relative weight between the type I and type II error rates. We let $k\in\{0.5,1,1.5\}$. A lower CER is more desirable. Top two rows in Figure 3 reports the results for the four designs. Across all designs, as the sample size increases, the rate at which CER declines, indicating potential diminishing return when sample size gets larger. In other words, the gain in reduction of error rates may lessen when sample size continues to increase. Considering the substantial costs associated with patient enrollment, these findings suggest finding a “sweet spot” of sample size that achieves a desirable tradeoff between cost and statistical properties. Comparing across designs, we find that except for the Standard SSE design, the performance of the other designs are similar for the same average sample size. Notably, the BESS SSR Cap design (design 2) seems to be in general the best across most cases as it shows slightly lower CER for fixed sample size. In Scenario 2, when $k=0.5$, the standard SSE performs well and is comparable to both the BESS SSR and BESS SSR Cap designs. In contrast, it is the losing design in the same scenario but with $k=1.5.$ This seems to suggest the standard SSE is a better design if the Frequentist type I error rate is of important consideration in design evaluation. However, in early-phase dose optimization trials, type I error rate is not the primary concern. For example, one would be much more concerned if a “GO” decision that recommend a wrong dose to further clinical development, or a “No Go” decision that fails to recommend a promising dose. These are measured by the FPR and FNR (See Appendix A.7 for detail) in the bottom two rows of Figure 3.